Examples of the Diagram Environment

Stolen from Various Sources

April 14, 1997

This ridiculous example shows how the package fits arrows in between the formulas, taking into account the exact size of every formula. (The box around the diagram shows how the shape of the entire diagram is made known to L^{A}TEX.) Note that diagonal arrows are fitted to either the tops or sides of the formulas, depending individual circumstances.

[figure: [ \begin{matrix}A_{00} & A_{01} \\ A_{10} & A_{11} \end{matrix} ]; B^{∗}; C; D; H^{2}(X, ω_{X} ⊗ L^{⊗(−n^{2}+n)}); a; c; u; b^{∗}; d; v; e]

(There are some additional diagrams at this point in the file, which you

can see if you add \usepackage{lamsarrow}\usepackage{pb-lams} at the

end of the list of included packages.)

These examples show how to simulate split arrows by placing the diagram on a finer grid than logically necessary.

[figure: A; B; C; D; α; β]

[figure: A; B^{∗}; C; D; E; a; c; u; b^{∗}; d; v; e]

Here are several “real life” examples from Bill Richter’s work:

[figure: K (Z/2, 8n + 1); K (Z/4, 8n − 1); E; ΣΩX ∧ ΩX; ΣΩX; X; K (Z/4, 8n).; Sq^{2}; Θ; π; H_{μ}; σ(α⊗α); σ; σ̃; α^{2}]

[figure: ΩΣA; Ω^{2} (ΣA ∧ ΣA); #; ΩB; F; Ω^{2} (B ∧ ΣA); A; X; B.; λ_{2}; δ; Ω(∂); h; π; J; Ω^{2}(∂∧id); e; f; E; h]

[figure: ΩS^{5}; (β); Ω (M_{2ι}^{5}); ΩS^{5}; G; J (S^{4} ∧ S^{4}); J_{2} (M_{2ι}^{4}); G_{2}; S^{8}; (α); M_{2ι}^{12}; S^{4}; M_{2ι}^{5}; S^{5}; Ω(pinch); δ; Ω(2ι); h_{2}; π; J; δ_{2}; ι; h_{2}; ι; E; τ̄; τ; ι; e; E; pinch]

Example by Anders Thorup (thorup@math.ku.dk), originally done with a package developed by himself and Steven Kleiman (kleiman@math.mit.edu):

[figure: H^{k}(B_{G} × N; Q) = H_{G}^{k}(N; Q); H_{G}^{k}(F_{j}; Q); H_{G}^{k}(M; Q); H^{k}(N; Q); H^{k}(F_{j}; Q); H^{k}(M; Q); f_{j}^{∗}; p^{∗}; f̃^{∗}; q_{j}^{∗}; i_{j}^{∗}; i^{∗}; f̃_{j}^{∗}=f_{j}^{∗}; f̃^{∗}=f^{∗}; i_{j}^{∗}]